OpenStudy (anonymous):

The perimeter of a standard sized rectangle is 36ft. The length is 2ft longer than the width. Find the dimensions. W=? ft L=? ft

OpenStudy (lgbasallote):

let W = width L = length L + L + W + W = perimeter L + L + W + W = 36 it also says the length is 2ft longer than width so L is 2ft longer than W L = 2 + W substitute that to L in that first equation L + L + W + W = 36 (2 + W) + (2 + W) + W + W = 36 2 + W + 2 + W +W +W = 36 4 + 4W = 36 4W = 36 - 4 4W = 32 W = 32/4 W = 8 can you solve for length now?

OpenStudy (anonymous):

10?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

formula for perimeter of rectangle is 2[length+width] i e p=2[l+w]. given parameters are: p=36, l=w+2 and w=w put the given parameters into the equation 2[w+2+w]=36 2[2w+2]=36 4w+4=36 4w=36-4 4w=32 w=32/4 w=8 therefore,width=8 and length=width+2=8+2=10.